OpenStudy (anonymous):

the subsitution property of equality allows you to say that if 7j +14k=28 and j =2 then ?

jimthompson5910 (jim_thompson5910):

well if j = 2, then you can replace every copy of j with 2 to get 7j +14k=28 7*2 +14k=28 14 + 14k = 28 now solve for k to get k = ??? (I'll let you solve for k)